calculate the bp of sol. containing 0.456 g of camphor dissloved in 31.4 g of acetone. Bp constant 56.30 C . If the molal elevation constant per 100g of acetone is 17.2

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
As from the given data:
Figure moles of camphor 
Molar mass = 152 g/mol 
No. Of mole = 0.456/152 = x 
Mass of CH3)2CO = 31.4 g 
Molality = x*1000/31.4 =m 
Rise of BP = k*m 
K= 17.2 for each 100g 
= 1.72 Per Kg 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Rise = 1.72 * x*1000/31.4 
BP = 56.30+ rise.